OpenStudy (anonymous):

please help figure out what to do next with my related rates prob. I've gotten to 5=400pi(dr/dt), what do I do next?

OpenStudy (anonymous):

anyone out there good with calculus?

OpenStudy (mertsj):

Solve for dr/dt. That's what you were asked to find...the rate at which the radius is changing.

OpenStudy (anonymous):

5/400 pi = dr/dt

OpenStudy (anonymous):

1/80 pi = dr/dt

OpenStudy (anonymous):

ok. that makes sense. thanks!